OpenStudy (anonymous):

suppose the population of a town is 100,000 in 1999. The population increased by 4.5% each year. what will be the population in 2005?

OpenStudy (anonymous):

u need to take out 5 % of 100000 for the first year... than take out 5% OF THE PREVIOUS YEAR ADDED WITH THE RISE IN POPULATION... DO THIS 5-6 TIMES UPTIL 2005

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

this is the formula A = P(1+r)^t to A = 100000(1+0.045)^6

OpenStudy (anonymous):

do u get it?

OpenStudy (anonymous):

255972?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ur number is going to be lower

OpenStudy (anonymous):

remember PEMDAS

OpenStudy (anonymous):

well thats what i got using that formula

OpenStudy (anonymous):

do u want to know the answer? first do parentheses did u do that?

OpenStudy (anonymous):

here hold on

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A = P(1+r)t (t is as an exponent) In the formula A = P(1+r)t, P is the principal\, r is the annual rate of interest, and A is the amount after t years. An account earning interest at a rate of 4% has a principal of $500,000. If no more deposits or withdrawls are made, about how much money will be in the account after five years? A $705,200 B $620,700 C $608,300 D $575,000 try this the answer is A = 500000(1.04)^5 A = 500000*1.2166529 A = $608,326.45, so guess it's c

OpenStudy (anonymous):

I based the formula on the compound interest formula which is Total amount = (Beginning amount) * (1 + rate)^years

OpenStudy (anonymous):

if u give up tell me..

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

here: Population = 100,000 *1.045^n (n is number of years since 1999. So, the population in 2005 (where n=6) is 130,226

OpenStudy (anonymous):

so the answer is right above this message